3 odd numbers between 1and 15 their sum is 30.you can repet the number
Answer by Edwin McCravy(20056)

Sorry, it can't be done.  3 ODD numbers always add up to an ODD number, never to
an EVEN number like 30.

Want proof?  Below, the computer has added every possible combination of three
odd numbers, repeated or not.  as you see, every one adds to an ODD number,
never to an EVEN number like 30.

1 + 1 + 1 = 3, which is an ODD number, NOT 30.
1 + 1 + 3 = 5, which is an ODD number, NOT 30.
1 + 1 + 5 = 7, which is an ODD number, NOT 30.
1 + 1 + 7 = 9, which is an ODD number, NOT 30.
1 + 1 + 9 = 11, which is an ODD number, NOT 30.
1 + 1 + 11 = 13, which is an ODD number, NOT 30.
1 + 1 + 13 = 15, which is an ODD number, NOT 30.
1 + 1 + 15 = 17, which is an ODD number, NOT 30.
1 + 3 + 3 = 7, which is an ODD number, NOT 30.
1 + 3 + 5 = 9, which is an ODD number, NOT 30.
1 + 3 + 7 = 11, which is an ODD number, NOT 30.
1 + 3 + 9 = 13, which is an ODD number, NOT 30.
1 + 3 + 11 = 15, which is an ODD number, NOT 30.
1 + 3 + 13 = 17, which is an ODD number, NOT 30.
1 + 3 + 15 = 19, which is an ODD number, NOT 30.
1 + 5 + 5 = 11, which is an ODD number, NOT 30.
1 + 5 + 7 = 13, which is an ODD number, NOT 30.
1 + 5 + 9 = 15, which is an ODD number, NOT 30.
1 + 5 + 11 = 17, which is an ODD number, NOT 30.
1 + 5 + 13 = 19, which is an ODD number, NOT 30.
1 + 5 + 15 = 21, which is an ODD number, NOT 30.
1 + 7 + 7 = 15, which is an ODD number, NOT 30.
1 + 7 + 9 = 17, which is an ODD number, NOT 30.
1 + 7 + 11 = 19, which is an ODD number, NOT 30.
1 + 7 + 13 = 21, which is an ODD number, NOT 30.
1 + 7 + 15 = 23, which is an ODD number, NOT 30.
1 + 9 + 9 = 19, which is an ODD number, NOT 30.
1 + 9 + 11 = 21, which is an ODD number, NOT 30.
1 + 9 + 13 = 23, which is an ODD number, NOT 30.
1 + 9 + 15 = 25, which is an ODD number, NOT 30.
1 + 11 + 11 = 23, which is an ODD number, NOT 30.
1 + 11 + 13 = 25, which is an ODD number, NOT 30.
1 + 11 + 15 = 27, which is an ODD number, NOT 30.
1 + 13 + 13 = 27, which is an ODD number, NOT 30.
1 + 13 + 15 = 29, which is an ODD number, NOT 30.
1 + 15 + 15 = 31, which is an ODD number, NOT 30.
3 + 3 + 3 = 9, which is an ODD number, NOT 30.
3 + 3 + 5 = 11, which is an ODD number, NOT 30.
3 + 3 + 7 = 13, which is an ODD number, NOT 30.
3 + 3 + 9 = 15, which is an ODD number, NOT 30.
3 + 3 + 11 = 17, which is an ODD number, NOT 30.
3 + 3 + 13 = 19, which is an ODD number, NOT 30.
3 + 3 + 15 = 21, which is an ODD number, NOT 30.
3 + 5 + 5 = 13, which is an ODD number, NOT 30.
3 + 5 + 7 = 15, which is an ODD number, NOT 30.
3 + 5 + 9 = 17, which is an ODD number, NOT 30.
3 + 5 + 11 = 19, which is an ODD number, NOT 30.
3 + 5 + 13 = 21, which is an ODD number, NOT 30.
3 + 5 + 15 = 23, which is an ODD number, NOT 30.
3 + 7 + 7 = 17, which is an ODD number, NOT 30.
3 + 7 + 9 = 19, which is an ODD number, NOT 30.
3 + 7 + 11 = 21, which is an ODD number, NOT 30.
3 + 7 + 13 = 23, which is an ODD number, NOT 30.
3 + 7 + 15 = 25, which is an ODD number, NOT 30.
3 + 9 + 9 = 21, which is an ODD number, NOT 30.
3 + 9 + 11 = 23, which is an ODD number, NOT 30.
3 + 9 + 13 = 25, which is an ODD number, NOT 30.
3 + 9 + 15 = 27, which is an ODD number, NOT 30.
3 + 11 + 11 = 25, which is an ODD number, NOT 30.
3 + 11 + 13 = 27, which is an ODD number, NOT 30.
3 + 11 + 15 = 29, which is an ODD number, NOT 30.
3 + 13 + 13 = 29, which is an ODD number, NOT 30.
3 + 13 + 15 = 31, which is an ODD number, NOT 30.
3 + 15 + 15 = 33, which is an ODD number, NOT 30.
5 + 5 + 5 = 15, which is an ODD number, NOT 30.
5 + 5 + 7 = 17, which is an ODD number, NOT 30.
5 + 5 + 9 = 19, which is an ODD number, NOT 30.
5 + 5 + 11 = 21, which is an ODD number, NOT 30.
5 + 5 + 13 = 23, which is an ODD number, NOT 30.
5 + 5 + 15 = 25, which is an ODD number, NOT 30.
5 + 7 + 7 = 19, which is an ODD number, NOT 30.
5 + 7 + 9 = 21, which is an ODD number, NOT 30.
5 + 7 + 11 = 23, which is an ODD number, NOT 30.
5 + 7 + 13 = 25, which is an ODD number, NOT 30.
5 + 7 + 15 = 27, which is an ODD number, NOT 30.
5 + 9 + 9 = 23, which is an ODD number, NOT 30.
5 + 9 + 11 = 25, which is an ODD number, NOT 30.
5 + 9 + 13 = 27, which is an ODD number, NOT 30.
5 + 9 + 15 = 29, which is an ODD number, NOT 30.
5 + 11 + 11 = 27, which is an ODD number, NOT 30.
5 + 11 + 13 = 29, which is an ODD number, NOT 30.
5 + 11 + 15 = 31, which is an ODD number, NOT 30.
5 + 13 + 13 = 31, which is an ODD number, NOT 30.
5 + 13 + 15 = 33, which is an ODD number, NOT 30.
5 + 15 + 15 = 35, which is an ODD number, NOT 30.
7 + 7 + 7 = 21, which is an ODD number, NOT 30.
7 + 7 + 9 = 23, which is an ODD number, NOT 30.
7 + 7 + 11 = 25, which is an ODD number, NOT 30.
7 + 7 + 13 = 27, which is an ODD number, NOT 30.
7 + 7 + 15 = 29, which is an ODD number, NOT 30.
7 + 9 + 9 = 25, which is an ODD number, NOT 30.
7 + 9 + 11 = 27, which is an ODD number, NOT 30.
7 + 9 + 13 = 29, which is an ODD number, NOT 30.
7 + 9 + 15 = 31, which is an ODD number, NOT 30.
7 + 11 + 11 = 29, which is an ODD number, NOT 30.
7 + 11 + 13 = 31, which is an ODD number, NOT 30.
7 + 11 + 15 = 33, which is an ODD number, NOT 30.
7 + 13 + 13 = 33, which is an ODD number, NOT 30.
7 + 13 + 15 = 35, which is an ODD number, NOT 30.
7 + 15 + 15 = 37, which is an ODD number, NOT 30.
9 + 9 + 9 = 27, which is an ODD number, NOT 30.
9 + 9 + 11 = 29, which is an ODD number, NOT 30.
9 + 9 + 13 = 31, which is an ODD number, NOT 30.
9 + 9 + 15 = 33, which is an ODD number, NOT 30.
9 + 11 + 11 = 31, which is an ODD number, NOT 30.
9 + 11 + 13 = 33, which is an ODD number, NOT 30.
9 + 11 + 15 = 35, which is an ODD number, NOT 30.
9 + 13 + 13 = 35, which is an ODD number, NOT 30.
9 + 13 + 15 = 37, which is an ODD number, NOT 30.
9 + 15 + 15 = 39, which is an ODD number, NOT 30.
11 + 11 + 11 = 33, which is an ODD number, NOT 30.
11 + 11 + 13 = 35, which is an ODD number, NOT 30.
11 + 11 + 15 = 37, which is an ODD number, NOT 30.
11 + 13 + 13 = 37, which is an ODD number, NOT 30.
11 + 13 + 15 = 39, which is an ODD number, NOT 30.
11 + 15 + 15 = 41, which is an ODD number, NOT 30.
13 + 13 + 13 = 39, which is an ODD number, NOT 30.
13 + 13 + 15 = 41, which is an ODD number, NOT 30.
13 + 15 + 15 = 43, which is an ODD number, NOT 30.
15 + 15 + 15 = 45, which is an ODD number, NOT 30.

Edwin